A 3.1g sample of unknown organic gas molecule compose of carbon, hydrogen and oxygen undergoes complete combustion to produce 4.4g CO2 and 2.7g H2O. Calculate the empirical and molecular formula of the gas if the molar mass is 62g/mol.


1
Expert's answer
2021-04-19T05:14:26-0400

Amounts of CO2 and H2O in moles are:


"n(CO_2)=\frac{4.4g}{44g/mol}=0.1mol"


"n(H_2O)=\frac{2.7g}{18g/mol}=0.15mol"


The mass of O2 can also be determined directly since the total mass of reactants is always equal to the total mass of products:

"m(O_2)=4.4+2.7-3.1=4.0g"

The amount of O2 in moles:


"n(O_2)=\frac{4.0g}{32g/mol}=0.125mol"


Therefore, the O2 : CO2 : H2O molar ratio is 0.125 : 0.1 : 0.15, which equals 5 : 4 : 6 as the lowest whole number ratio.

Now the overall equation can be written:

nCxHyOz + 5O2 --> 4CO2 + 6H2O

For the equation to become balanced, there should be 4 carbon atoms, 12 hydrogen atoms, and 4 oxygen atoms on the left side, in other words, x : y : z = 4 : 12 : 4. The lowest whole number ratio of this (1 : 3 : 1) determines the empirical formula, which is CH3O. To calculate the molecular formula, the molar masses of the empirical formula unit and the unknown compound should be compared:

"M(CH_3O)=12+3\times1+16=31g/mol" , which is 2 times less than the actual molar mass (62 g/mol). Therefore, the subscripts of the empirical formula should be multiplied by 2 giving C2H6O2.


Answer:

Empirical formula CH3O

Molecular formula C2H6O2